OpenStudy (anonymous):

what is the point-slope form equation for the line that passes through the point (-2,1) and has a slope of -3

Nnesha (nnesha):

\[\huge\rm y-y_1 = m (x-x_1)\] point slope form where m is slope x_1 and y _1 is the order pair (-2,1)

OpenStudy (anonymous):

what would be the answer with the numbers put in to the formula.

Nnesha (nnesha):

replace variable by their values for example m = -3 so \[\huge\rm y-y_1 = -3(x-x_1)\] now replace x_1 and y_1 by their values

OpenStudy (anonymous):

Can you please just put the numbers into the formula so I can understand better.

Nnesha (nnesha):

for example if order pair is (4 , -5) i will plug in x_1 = 4 y_1 = -5 \[\huge\rm y-(\color{reD}{-5}) = -3(x-\color{red}{4})\] then i have to distribute negative 1 by negative 5 so final answer would be \[\huge\rm y+\color{reD}{5} = -3(x-\color{red}{4})\] y+5 negative times negative = positive

Nnesha (nnesha):

now its ur turn